# LLMs Can’t Handle Peer Pressure: Crumbling under Multi-Agent Social Interactions

### v2

@@ Abstract
Large language models (LLMs) are increasingly deployed in multi-agent systems (MAS) as components of collaborative intelligence, where peer interactions dynamically shape individual decision-making. Although prior work has focused on conformity bias, we extend the analysis to examine how LLMs form trust from previous impressions, resist misinformation, and integrate peer input during interaction, key factors for achieving collective intelligence under complex social dynamics. We present Kairos, a benchmark simulating quiz contests with peer agents of varying reliability, offering fine-grained control over conditions such as expert–novice roles, noisy crowds, and adversarial peers. LLMs receive both historical interactions and current peer responses, allowing systematic investigation into how trust, peer action, and self-confidence influence decisions. As for mitigation strategies, we evaluate prompting, supervised fine-tuning, and reinforcement learning—Group Relative Policy Optimization (GRPO)—across multiple models. Our results reveal that GRPO with multi-agent context combined with outcome-based rewards and unconstrained reasoning achieves the best overall performance, but also decreases the robustness to social influence compared to Base models. The code and datasets are available at: https://github.com/declare-lab/KAIROS.
@@ 1 Introduction
Large Language Models (LLMs) are increasingly integrated into multi-agent systems (MAS), where they must interact, reason, and collaborate with other agents (Chen et al. 2024b; Tran et al. 2025). However, like humans, LLMs are vulnerable to social and cognitive biases such as conformity, overconfidence, and herd behaviour (Piatti et al. 2024; Weng, Chen, and Wang 2025; Yan et al. 2025). When exposed to peer responses, LLMs may adjust their outputs not only to align with group consensus but also due to misplaced trust in unreliable agents (Cho, Guntuku, and Ungar 2025). These tendencies pose a critical challenge in collective decision-making, where a single flawed response can propagate across agents—cascading through the system—and ultimately compromise the reliability of the entire multi-agent framework.
@@
While previous studies have explored conformity in isolated settings (Weng, Chen, and Wang 2025; Zhu et al. 2024), a comprehensive framework for simulating interactive social environments and systematically evaluating LLM behaviour under varying conditions of trust, peer influence, and self-confidence is still lacking. To address this, we introduce Kairos, a benchmark for assessing LLMs in socially grounded, multi-agent scenarios. It simulates quiz-style multiple-choice contests where the model interacts with peer agents of varying reliability, relying on both historical trust and self-confidence to make decisions. We evaluate model behaviour using four metrics: accuracy, the overall task success rate; utility, the ability to correct errors through peer input; resistance, the ability to maintain its stance when correct; and robustness, the change in accuracy between original and peer influence settings, reflecting stability under social interactions.
@@
Beyond measuring susceptibility to social cues, our goal is to develop and evaluate mitigation strategies that enhance model performance within multi-agent social simulations. We explore three main approaches: prompting, supervised fine-tuning, and reinforcement learning (GRPO). Within GRPO, we systematically vary four dimensions: system prompt design, reward formulation, inclusion of multi-agent context, and data filtering strategies. Our experiments show that GRPO, when trained under carefully designed MAS conditions with outcome rewards, significantly outperforms prompting and SFT baselines, improving original task performance while maintaining robustness to social perturbations. In contrast, other strategies, despite boosting surface-level accuracy, often fail to generalise under protocol shifts, revealing a persistent fragility in social reasoning. These findings underscore that improving accuracy alone is insufficient and that robust reasoning under social interference remains a key bottleneck in multi-agent generalisation.
Our work aims to systematically evaluate and improve how LLMs perform in socially interactive, trust-sensitive environments. To that end, we make the following key contributions:
@@
  • A novel social interaction benchmark: We introduce a quiz-style multi-agent simulation that includes a controlled variation of peer reliability, historical trust level, and self-confidence influence, allowing us to measure how LLMs adapt to complex social cues.
@@
  • Comprehensive analysis of social behaviours in LLMs: We evaluate model behaviours across architectures and training regimes, identifying patterns of interaction dynamics, trust sensitivity, and reward alignment in multi-agent settings.
@@
  • Evaluation of training and prompting strategies: We compare prompting-based, supervised fine-tuning (SFT), and reinforcement learning (via GRPO) methods under different objectives. Our results show that GRPO with outcome reward and MAS context yields the most robust performance.

### v3

@@ Abstract
Large language models (LLMs) are increasingly integrated into multi-agent systems (MAS), where peer interactions shape individual decisions. While prior work has mainly examined conformity bias, we broaden the view to include how LLMs build rapport from prior interactions, discern and integrate high-quality peer information, and resist misleading inputs—abilities essential for achieving collective intelligence under complex social dynamics. We introduce Kairos, a benchmark that simulates quiz-style collaboration with peer agents whose rapport levels and behaviors can be precisely controlled in both historical interactions and the current round. This unified setup enables systematic analysis of how rapport, peer actions, and the model’s self-confidence jointly influence decision-making. Using Kairos, we evaluate prompting, supervised fine-tuning, and reinforcement learning via Group Relative Policy Optimisation (GRPO). Results show that model scale is a primary factor moderating susceptibility to social influence: larger models are more resilient and benefit from prompting-based mitigation, whereas smaller models remain vulnerable. Only carefully configured GRPO training yields consistent robustness and performance gains for small models.
@@ 1 Introduction
Large Language Models (LLMs) are increasingly integrated into multi-agent systems (MAS), where they must interact, reason, and collaborate with other agents chen2024survey; tran2025multi. However, like humans, LLMs are vulnerable to social and cognitive biases such as conformity, overconfidence, and herd behaviour piatti2024cooperate; wengwe; yan2025beyond. When exposed to peer responses, LLMs may adjust their outputs not only to align with group consensus but also due to misplaced trust in unreliable agents cho2025herd. These tendencies pose a critical challenge in collective decision-making, where a single flawed response can propagate across agents, cascading through the system, and ultimately compromise the reliability of the entire multi-agent framework.
@@
While prior studies have examined conformity in controlled, isolated settings wengwe; zhu2024conformity, the field still lacks a comprehensive framework for simulating interactive social environments and systematically evaluating LLM behaviour under varying conditions of rapport, current peer behaviour, and self-confidence. To fill this gap, we introduce Kairos, a benchmark designed to assess LLMs in socially grounded, multi-agent scenarios. It simulates quiz-style multiple-choice contests in which the model engages with peer agents of varying reliability, leveraging both historical rapport and its own confidence to make decisions. Model behaviour is evaluated using four metrics: accuracy, the overall task success rate; utility, the ability to correct its own errors with peer input; resistance, the ability to maintain correct judgments despite misleading peers; and robustness, the change in accuracy from the original to the socially influenced setting, capturing stability under social interaction.
@@
Beyond measuring susceptibility to social cues, our aim is to identify mitigation strategies that enhance model performance in multi-agent social simulations. We investigate three approaches: prompting, supervised fine-tuning, and reinforcement learning via GRPO. For GRPO, we systematically vary four factors: system prompt design, reward structure, incorporation of multi-agent context, and data filtering strategies. Our results show that GRPO, when trained with carefully configured MAS signals and outcome-based rewards, substantially outperforms prompting and SFT baselines, improving core task performance while preserving robustness against social perturbations. In contrast, alternative methods, though sometimes improving surface-level accuracy, often fail to generalise across protocol shifts, revealing persistent weaknesses in social reasoning. These findings highlight that accuracy alone is insufficient; resilient reasoning under social interference remains a central challenge for multi-agent generalisation.
Our work provides a systematic evaluation of LLM performance in socially interactive, trust-sensitive settings. Our key contributions are:
@@
  • A novel social-interaction benchmark: We introduce a quiz-style multi-agent simulation that manipulates peer reliability, historical rapport, and self-confidence, enabling precise measurement of how LLMs respond to complex social cues.
@@
  • A comprehensive study of LLM social behaviour: We analyse model behaviours across architectures and training regimes, revealing patterns in interaction dynamics, trust sensitivity, and reward alignment in multi-agent environments.
@@
  • A comparative evaluation of mitigation strategies: We examine prompting, supervised fine-tuning, and GRPO-based reinforcement learning under different configurations, finding that GRPO with outcome rewards and MAS context achieves the most robust gains.
We present a multi-agent interactive benchmark, Kairos, designed to simulate socially grounded scenarios and assess LLM behaviour within them. Unlike benchmarks focused solely on conformity wengwe, Kairos targets how LLMs interpret, utilise, or resist signals from other agents, even when they are clearly unreliable, based on perceived reliability, current social context, and self-confidence. We begin by describing the details of collecting dynamic evaluation data in Section 2.1 and then explaining the corresponding metrics in Section 2.2.
2.1 Data Collection
A detailed breakdown of the training and evaluation datasets is presented in Figure 1, covering the four domains of Reasoning, Knowledge, Social, and Creativity. The datasets are carefully partitioned to ensure strict non-overlap and a clear distributional shift between training and evaluation. Additional details are provided in appendix A.
Figure 1: Left: Training dataset (N=10,000). Right: Test dataset (N=3,000). The inner ring groups tasks by category — Training: Reasoning 37.5%, Knowledge 21.1%, Social 20.8%, Creativity 20.5%; Test: Reasoning 33.3%, Knowledge 22.3%, Social 22.2%, Creativity 22.2%. The outer ring breaks each category into individual datasets; wedge labels give original instance counts.
Evaluation Dataset Source.
To evaluate models under realistic social dynamics, we draw benchmark questions from a diverse set of established datasets spanning four domains—Reasoning, Knowledge, Social, and Creativity. The Reasoning domain combines logic reasoning wengwe, filtered BIG-Bench Hard bbh, code execution tasks from LiveCodeBench jain2024livecodebench, and level 4–5 problems from MATH-500 lightman2023lets. The Knowledge domain uses TruthfulQA lin2021truthfulqa and MMLU-Pro wang2024mmlu to evaluate factual and broad-domain knowledge. The Social domain includes CommonsenseQA 2.0 talmor2022commonsenseqa and Social IQ sap2019socialiqacommonsensereasoningsocial for intuitive and socially grounded reasoning. Finally, the Creativity domain incorporates MacGyver tian2023macgyver for situational problem-solving and BrainTeaser jiang2023brainteaser for lateral-thinking challenges.
All tasks are reformulated into a unified multiple-choice question answering (MCQA) format. We adopt MCQA for its straightforward answer extraction and deterministic evaluation, in contrast to open-ended formats that require LLM-based judging and introduce additional uncertainty (see Section F.1 for details). For originally open-ended datasets (e.g., LiveCodeBench, MATH-500), we generate plausible distractors using Llama3.1-8B and validate them through automated checks and human review. The MacGyver dataset is converted into a binary solvability judgment while retaining its creative reasoning characteristics.
Dynamic Evaluation Dataset Construction.
To evaluate a model’s robustness in socially rich settings, we first infer its underlying beliefs—its preferred answer and corresponding confidence for each benchmark question. This allows us to construct tailored evaluation scenarios that probe the model’s own epistemic commitments rather than relying on a fixed benchmark. As a result, Kairos is dynamically instantiated for each model, adapting to its responses and beliefs to stress-test its social reasoning.
Step 1: Extracting the Model’s Original Beliefs
We begin by presenting the model with the original benchmark question and recording its direct output. From this output, we extract the discrete final answer label (e.g., “A”, “B”, “C”, or “D”), which we refer to as the model’s original answer and treat as its stated belief.
To estimate the model’s confidence in this belief, we adopt a sampling-based uncertainty estimation procedure similar to Self-Consistency. For each input $\mathbf{x}$, we generate $T$ full solutions using stochastic decoding and extract the final answer label $y_{t}\in\{1,\dots,K\}$ from each generation:
$\{y_{t}\}_{t=1}^{T}.$
These samples induce an empirical predictive distribution over answer options,
$\bar{p}_{k}=\hat{p}(y=k\mid\mathbf{x})=\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}[y_{t}=k],\quad k=1,\dots,K,$
collectively denoted as $\bar{\mathbf{p}}$. Confidence is quantified via the predictive entropy
$\mathcal{H}[\bar{\mathbf{p}}]=-\sum_{k=1}^{K}\bar{p}_{k}\log\bar{p}_{k}.$
The model’s final belief is taken to be the majority answer—the option with the highest empirical probability under $\bar{\mathbf{p}}$. Finally, samples are categorized into high- or low-confidence by comparing their entropy to the dataset-wide median: those above the median are labeled low-confidence and those below are labeled high-confidence.
Step 2: Simulating Social Scenarios with Targeted Interventions
Once the model’s belief, its predicted answer, and the associated confidence have been established, we construct a multi-agent simulation to examine how it responds to social influence. Each simulation consists of two components: simulated interaction history and current question round.
The interaction history mimics prior rounds of a quiz-style contest. For each past round, we provide the previous questions, the model’s own answers, and the responses of peer agents. This context allows us to simulate the buildup of agent-specific rapport based on how consistently each peer has aligned with the evaluated model’s prior answers.
In the current round, the model is given a new question along with responses from peer agents that are intentionally crafted to either align with or challenge its previously expressed belief. This alignment is defined along three behavioural modes: support, oppose-hard, and oppose-easy. If the model’s original answer is correct, support agents reinforce it by repeating the correct answer, oppose-hard agents challenge it by selecting the most plausible incorrect option, and oppose-easy agents offer minimal resistance by selecting the least plausible incorrect answer. Conversely, if the model’s answer is incorrect, support agents echo the same incorrect response, oppose-hard agents provide a different but highly plausible incorrect alternative, and oppose-easy agents present the correct answer. This targeted construction allows us to simulate varying degrees of social pressure and systematically assess how the model responds to both reinforcing and contradicting social signals.
Taken together, this setup allows Kairos to probe three key axes of a model’s social sensitivity:
  • Peer Agent Rapport Level: By manipulating the degree to which past peer answers agree with the model’s own history (ranging from $0\%,25\%,50\%,75\%,100\%$}), Kairos evaluates how perceived reliability and social closeness modulate the model’s trust in others.
  • Peer Agent Behaviour in the Current Round: Because peer responses are constructed directly from the model’s belief distribution in Step 1, their support, oppose-hard, or oppose-easy behaviours constitute a model-tailored stress test. This allows us to examine how the model handles aligned versus adversarial signals conditioned on its own epistemic commitments.
  • Self-Belief Strength: By conditioning interventions on the model’s estimated confidence (high vs. low entropy), Kairos tests whether self-belief moderates susceptibility to social influence—i.e., whether uncertainty increases conformity and high confidence supports resistance.
This formulation enables a structured examination of how peer agreement, perceived trustworthiness, and the model’s own self-belief jointly shape behaviour under social pressure. Because peer responses are derived from each model’s belief distribution, every model encounters a tailored instantiation of Kairos. To enable cross-model comparison despite differing baselines, we propose normalized robustness metrics that quantify behavioural shifts under social interaction.
2.2 Evaluation Metrics
Accuracy and Robustness.
Utility & Resistance.
[…]
We define two complementary measures of model $M$’s performance under the Kairos evaluation. The utility $U_{M}$ quantifies the fraction of instances that were originally incorrect but become correct under Kairos, while the resistance $R_{M}$ captures the fraction of instances that were originally correct and remain correct. Formally,
$U_{M}=\frac{\sum_{i=1}^{N}\mathbf{1}\{x_{i}=0\wedge y_{i}=1\}}{\sum_{i=1}^{N}\mathbf{1}\{x_{i}=0\}}\;\in[0,1],\qquad R_{M}=\frac{\sum_{i=1}^{N}\mathbf{1}\{x_{i}=1\wedge y_{i}=1\}}{\sum_{i=1}^{N}\mathbf{1}\{x_{i}=1\}}\;\in[0,1].$
2.3 Why Robustness Matters in Multi-Agent Settings
While accuracy improvements are important, robustness—the stability of performance between the Original and Kairos evaluations—is equally crucial. A widening gap between the two signals heightened sensitivity to social context. This fragility raises several concerns:
[…]
  • Vulnerability to Social Pressure: Models generally lose more correct predictions than they gain from peer corrections, suggesting they are more easily swayed into errors than guided toward better answers. This mirrors human social pitfalls and enables mistake propagation across agents.
  • Hidden Brittleness: Reinforcement learning often increases surface-level accuracy while widening the Original–Kairos gap, revealing overfitting to evaluation formats. Without robustness, accuracy gains fail to generalise under distributional or social shifts.
  • System-Level Risk: In collective settings, a socially induced error by one agent can cascade through the group, amplifying misinformation and destabilising shared outputs. Such susceptibilities also open new avenues for adversarial manipulation.
Robustness is therefore not secondary to accuracy but a fundamental requirement for deploying LLMs in multi-agent environments. Improving accuracy without ensuring robustness risks creating systems that perform well in controlled benchmarks yet falter in realistic, socially entangled conditions.
3 Experimental Setup
3.1 Mitigation Strategies
To improve the robustness of LLMs in socially interactive environments, we examine three categories of mitigation strategies—Prompting, Supervised Fine-Tuning (SFT), and Reinforcement Learning via Group Relative Policy Optimisation (GRPO)—each designed to strengthen the model’s ability to reason accurately while managing peer influence in multi-agent settings. For clarity, all comparisons are made against the same base model per family, and every SFT or GRPO variant within that family is fine-tuned from this same Base checkpoint.
3.1.1 Prompting
Following wengwe, we also explore two different prompting strategies: Empowered and Reflective.
3.1.2 Supervised Fine-Tuning (SFT)
3.1.3 Reinforcement Learning
  1. 1.
    Multi-Agent vs. Non-Multi-Agent Context: In the MAS configuration, training inputs include the full history of prior questions and peer agent responses, teaching the model to navigate social interference while improving task performance. In contrast, the Non-MAS setting removes all social context, serving as a control setup to test whether simply improving a model’s competence reduces its susceptibility to peer influence.
  2. 2.
  3. 3.
  4. 4.
3.2 Training Dataset Construction for Mitigation Strategies
To avoid contamination and enforce a strict train–test separation, we construct the training set from disjoint sources spanning the same domains as Kairos: reasoning, knowledge, social, & creativity.
Reasoning: Includes BBH bbh and LiveCodeBench jain2024livecodebench—both used under a strict train–test split to avoid data leakage—along with MathQA amini-etal-2019-mathqa for math word problems and Winogrande ai2winogrande for commonsense reasoning. Knowledge: Uses PopQA mallen2023llmmemorization and SimpleQA wei2024measuringshortformfactualitylarge for factual question answering. Social: Combines Social yuan2024measuring and Moral Stories Emelin2021MoralSS to promote understanding of social norms and moral reasoning. Creativity: Incorporates ProtoQA ProtoQA for prototypical reasoning and RiddleSense for metaphorical comprehension.
We process all datasets in the same way as Kairos. For each interaction history, we ensure a balanced mix of round types (both support and oppose). In open-ended datasets such as SimpleQA, distractor answers are generated automatically. More details can be found in appendix A.
4 Results
4.1 Overall Results
Table 1 reports results for 11 models evaluated on Kairos across different prompting strategies. We observe distinct behaviors between smaller ($\leq$ 32B) and larger (>> 32B) models. Under empowered prompting, smaller models exhibit performance gains in both Original ($\uparrow 4.37\%$) and Kairos ($\uparrow 0.95\%$) settings. However, because the improvement in isolated performance outpaces the gains in social contexts, this strategy paradoxically exacerbates the robustness gap ($O-K\Delta$). This suggests that while empowerment helps elicit latent capabilities in smaller models, it does not proportionally enhance resilience to social interference.
In contrast, larger models show saturation in Original accuracy under base prompting, leaving limited room for improvement ($\downarrow 0.08\%$); consequently, empowerment primarily boosts Kairos accuracy ($\uparrow 3.10\%$), effectively closing the robustness gap and resulting in a positive $\Delta$. Finally, reflected prompting proves detrimental for smaller models, degrading Kairos performance ($\downarrow 2.83\%$) and widening the robustness gap compared to the base setting, likely due to hallucinations or confusion induced during the self-reflection process. Larger models remain relatively stable under reflection but still underperform compared to the empowered setting.
Models Base Empowered Reflected
Original ($\uparrow$) Kairos ($\uparrow$) O–K $\Delta$ ($\uparrow$) Original ($\uparrow$) Kairos ($\uparrow$) O–K $\Delta$ ($\uparrow$) Original ($\uparrow$) Kairos ($\uparrow$) O–K $\Delta$ ($\uparrow$)
Qwen2.5-3B 47.93% 48.77% $\mathbf{+2.4\%}$ 56.06% 47.87% -14.6% 47.93% 47.27% -1.4%
Qwen2.5-7B 58.50% 52.27% -10.0% 65.74% 54.07% $\mathbf{-17.8\%}$ 58.50% 55.33% -5.4%
Qwen2.5-14B 64.00% 58.43% -8.7% 68.23% 62.50% -8.4% 64.00% 59.19% -7.5%
Llama3.2-3B 47.90% 43.81% -7.8% 48.43% 44.70% -7.7% 47.90% 38.40% -19.8%
Llama3.1-8B 56.50% 52.54% -7.0% 61.03% 53.04% -13.1% 56.50% 40.59% $\mathbf{-28.1\%}$
Llama3.3-70B 67.97% 68.17% +0.3% 68.47% 69.60% +1.6% 67.97% 66.80% -1.7%
QWen2.5-32B 69.30% 67.37% -2.8% 70.90% 66.73% -5.9% 69.30% 65.43% -5.6%
QWen2.5-72B 69.33% 69.43% +0.1% 69.23% 71.07% $\mathbf{+2.7\%}$ 69.33% 68.73% -0.9%
GPT-OSS 120B 86.67% 80.87% -6.7% 87.20% 83.97% -3.7% 86.67% 85.47% -1.4%
Gemini-2.5-Pro 89.33% 79.93% -10.5% 88.23% 88.17% -0.1% 89.33% 87.50% -2.0%
GPT-5 90.17% 88.90% -1.4% 89.90% 90.00% +0.1% 90.17% 90.03% $\mathbf{-0.1\%}$
Avg (LLMs $\leq$ 32B) 57.36% 53.87% -5.65% 61.73% 54.82% -11.25% 57.36% 51.04% -11.30%
Avg (LLMs >> 32B) 80.69% 77.46% -3.64% 80.61% 80.56% +0.12% 80.89% 79.71% -1.22%
Table 1: Evaluation of model robustness under Kairos. The table summarises Original and Kairos accuracies and their relative $\text{O--K}~\Delta$ (percentage change) across multiple model families, sizes, over prompting strategies. The maximum and minimum $\text{O--K}~\Delta$ values are highlighted in bold.
4.2 Detailed Results
Beyond prompting, we also examine training-based mitigation strategies for improving robustness to social interference. Due to computational constraints, we limit SFT and GRPO experiments to models under 32B parameters. Detailed dataset-level results are provided in Table 811, while Table 2 presents a consolidated comparison.
Type Qwen2.5-3B Qwen2.5-7B Qwen2.5-14B LLama3.2-3B LLama3.1-8B
Original Kairos O–K $\Delta$ Original Kairos O–K $\Delta$ Original Kairos O–K $\Delta$ Original Kairos O–K $\Delta$ Original Kairos O–K $\Delta$
Base 47.9 48.8 1.8 58.5 52.3 -10.6 64.0 58.4 -8.7 47.9 43.8 -8.6 56.5 52.5 -7.0
Empowered 56.1 47.9 -14.6 65.7 54.1 -17.7 68.2 62.5 -8.4 48.4 44.7 -7.7 61.0 53.0 -13.1
Reflected 47.9 47.3 -1.4 58.5 55.3 -5.4 64.0 59.2 -7.5 47.9 38.4 -19.8 56.5 40.6 -28.1
SFT 50.1 46.9 -6.5 56.7 44.0 -22.4 65.3 48.8 -25.3 45.0 39.4 -12.6 49.3 42.1 -14.6
GRPO-MAS-DS-DR 54.8 51.7 -5.7 66.6 62.0 -6.9 75.6 69.5 -8.0 51.1 46.1 -9.8 60.4 55.7 -7.9
GRPO-MAS-DS-DR-LConf 52.5 48.8 -7.0 63.4 54.3 -14.4 70.1 60.9 -13.2 51.7 44.7 -13.5 58.7 52.3 -10.9
GRPO-MAS-DS-DR-LCorr 55.6 47.5 -14.6 63.3 49.9 -21.2 68.6 45.8 -33.3 52.0 45.9 -11.7 60.8 47.6 -21.6
GRPO-MAS-DS-OR 57.4 52.8 -7.9 67.4 62.5 -7.2 73.3 70.3 -4.1 52.0 48.3 -7.2 58.3 56.4 -3.3
GRPO-MAS-NS-OR 61.7 57.9 -6.1 70.3 65.5 -6.8 76.4 71.5 -6.5 55.7 51.3 -8.0 63.8 57.3 -10.2
GRPO-nonMAS-DS-DR 57.6 51.3 -11.0 64.5 59.3 -8.0 72.8 62.5 -14.1 55.5 45.0 -19.0 59.3 49.1 -17.2
GRPO-nonMAS-DS-OR 56.3 50.8 -9.7 68.6 56.4 -17.8 71.1 65.8 -7.4 55.5 44.2 -20.4 55.9 51.6 -7.7
GRPO-nonMAS-NS-OR 62.7 53.8 -14.2 72.7 57.7 -20.7 77.5 65.5 -15.6 58.2 50.2 -13.6 63.8 56.1 -12.0
Table 2: Comparison of Original accuracy, Kairos accuracy, and O–K $\Delta$ across different models and mitigation configurations. For each model family, all SFT and GRPO variants are fine-tuned from the same Base checkpoint, enabling consistent comparison of how prompting, SFT, and GRPO influence robustness under social interaction.
GRPO Boosts Accuracy But Not Robustness.
As shown in Table 2, training with GRPO consistently improves both Original and Kairos performance compared to Supervised Fine-Tuning (SFT). On average, GRPO yields a +12.3% gain in Original accuracy and a +16.4% gain in Kairos accuracy, confirming that reinforcement learning enhances models’ ability to operate in socially interactive settings where passive imitation alone is insufficient.
These gains, however, are not uniform across robustness. While many GRPO-trained models achieve higher accuracy at the expense of robustness, some variants, particularly those trained with MAS context, also yield robustness improvements. This indicates that GRPO enhances competence overall, but its effect on resilience to social pressure depends on how the training context is structured.
MAS Context Enhances Both Performance and Robustness.
NS-OR Configuration Yields Best Accuracy-Robustness Trade-off.
Filtering by Confidence Helps Accuracy but Harms Robustness.
Key Takeaways GRPO improves accuracy over SFT, but can reduce robustness in certain settings. Incorporating social context (MAS) during GRPO training boosts accuracy and, for larger models, enhances robustness, while smaller models may experience a drop. Debate-style reasoning enforced through system prompts or rewards shows no improvements in robustness or accuracy, suggesting models benefit more from simpler objectives.
4.3 Analysis on Transition Effect and Peer Impression
We analyze how LLMs behave under multi-agent social context, focusing on two phenomena: (1) how models transition between correct and incorrect answers when exposed to peer input, and (2) how prior rapport with peers shapes the strength of social influence. A detailed breakdown is provided in Appendix D.
Models Lose More Correct Answers Than They Gain.
Across models and training regimes, we observe a consistent net accuracy drop when moving from the initial context to the multi-agent setting. This arises because losses from incorrect transitions (correct $\rightarrow$ wrong) outweigh the gains from utility transitions (wrong $\rightarrow$ correct). Resistance transitions (correct $\rightarrow$ correct) dominate overall behaviour—accounting for roughly 65% of all transitions—indicating a structural bias toward preserving initial judgments rather than revising them. Training modulates this balance: SFT increases error correction but reduces resistance confidence, while GRPO restores strong resistance but suppresses utility confidence (e.g., Qwen-14B utility drops from 0.737 to 0.167). Larger models generally show stronger resistance and utility than smaller ones, though training narrows these scale-driven advantages. Despite these improvements, models still struggle to confidently and systematically revise erroneous beliefs based on peer signals.
Rapport Modulates Peer Influence.
Prior rapport systematically alters how models respond to current peer behavior. Higher rapport increases resistance in support settings—making models more likely to maintain a correct answer when peers agree—but decreases resistance in oppose-hard and oppose-easy, where high rapport amplifies conformity to incorrect peers. Utility transitions show the opposite trend, indicating that rapport affects action selection rather than confidence. Across model families, this produces a persistent asymmetry: models over-trust supportive peers and underperform when required to reject misleading consensus, with an average 31.7-point resistance gap between support and oppose-hard. Training further shapes these dynamics: GRPO strengthens both resistance and utility under challenging opposition, whereas SFT improves behavior in easier settings but degrades performance in harder corrective scenarios. Overall, rapport acts as a powerful modulator of social influence, amplifying agreement tendencies and revealing a systematic vulnerability to misleading group consensus.
4.4 Additional Analysis
We provide supplementary analyses in Appendix F, covering the evaluation differences between MCQ and open-ended formats, the impact of historical rapport on model behaviour, and how the model’s confidence on history questions shapes its current decisions.